OpenStudy (anonymous):

How to solve this? lim x→pi/2 (cos x/ 1- 4x^2/pi^2)

OpenStudy (anonymous):

The question is actually like this:

OpenStudy (anonymous):

Use L-Hospital maybe? -\[\Large \frac{-\sin x} {\frac{8x}{\pi^2}}=\lim_{x \rightarrow \frac{\pi}{2}} \frac{-\pi ^2 \sin x}{8x}\] put pi/2 now

OpenStudy (anonymous):

The answer is -1/2. I just need to see the steps or method on solving this

OpenStudy (anonymous):

do u know l hospital?

OpenStudy (anonymous):

yeah why